OpenStudy (anonymous):

If y = 3x – 18 and x = 3, what is the value of y?

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

Because y=3x3-18 y=9-18 y=-9

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome :)